OpenStudy (anonymous):

The base of a trianlge exceeds the height by six centimeters. If the area is 36 centimeters, find the length of the base and the height of the traingle???

OpenStudy (anonymous):

\[A=\frac{1}{2}bh\] \[A=\frac{1}{2}h(h+6)\]

OpenStudy (anonymous):

because b=h+6

OpenStudy (anonymous):

(1/2)( x)(x+6) = 36 x^2 +6x = 72 x^2 +6x -72=0 (x-6)(x+12)=0 x=6 or -12 (rej.)

OpenStudy (anonymous):

height= 6, base= 12

OpenStudy (anonymous):

thank you!!!!!!!